# STrack: A Reliable Multipath Transport for AI/ML Clusters

### v1

@@ 3.3. STrack Reliability
upon the arrival of each packet, a receiver sends an Selective Acknowledgment (SACK) packet if one of the following conditions are met: (1) a sufficient number of bytes have been received; (2) a packet with current expected Packet Sequence Number (EPSN) is received; (3) a probing packet is received; (4) a packet that requests for ACK is received.
@@
The receiver records the lowest PSN since the last SACK. The next SACK is sent with the bitmap segment that is associated with that lowest PSN. This is because the packet retransmission from the lowest unacknowledged PSN and thus, we need the latest arrival states for the lowest PSN.
@@
Loss is inferred if the out-of-order packets counter carried at SACK exceeds a threshold. The threshold is recommended to be set to the maximum of the current congestion window’s worth of packet number and min_threshold (say 5). This constant threshold prevents spurious retransmissions when the congestion window is less than min_threshold packets. The rationale for using a threshold based on the congestion window is that if a packet does not get acknowledged within a whole window’s time, it is safe to consider it as lost. This also helps us avoid the bitmap size at receiver/sender
@@ 4.1. Experiment Setup
We compare STrack with RoCEv2 (rocev2, ), which is widely deployed transport over Ethernet in AI/ML network. We configure the constants for STrack congestion control as suggested in  Table 1 and set the maximum number of paths as $256$ for adaptive load balancing and switch ECN mark threshold $K_{min}=25\%BDP$ and $K_{max}=75\%BDP$. Switch is operated without any advanced features enabled, and drops packets if queuing exceeds $5$ BDP. RoCEv2 utilizes DCQCN (dcqcn, ) as its primary congestion control mechanism and supports single-path communication. Additionally, it necessitates Priority Flow Control (PFC) to be enabled throughout the entire network to prevent packet loss caused by buffer overflow. We implemented shared buffer architecture for PFC to best demonstrate the performance of RoCEv2. We configured the total switch buffer size to be $256$MB for a total capacity of $51.2$Tbps and scaled the buffer size accordingly for switches with different radices. For instance, the switch buffer size is set to 128MB for a 25.6Tbps switch. We set the ECN threshold to one BDP for DCQCN and configure Lossless for RoCEv2.
@@ 4.2. Synthetic Traffic
Figure 11 samples the switch queuing delay as the simulation time progresses in the $16MB$ experiment in Figure 11 for STrack and RoCEv2. There are a total of $16K$ switch queues, $8$K queues from ToR to Spine Switch and the other $8$K queues from Spine to ToR. To avoid too many logs to be processed, we only generate logs if the queuing delay is larger than $8$us (the base RTT). Hence, if we do not see any points at one time moment in Figure 11, it means that across the $16$K queues, there is not any queue whose queue depth exceeds the BDP at that time point. Figure 11 shows that before $370$us, STrack’s adaptive packet spraying experiences hash collision, as it is trying to skip the bad paths, and find the good paths. After $370$us, all the flows get the settle down to the good paths, and thus, the switch queuing delays do not exceed $8$us after $370$us. In contrast, RoCEv2 continues experiencing hash collisions, which leads to the queuing delay up to $250$us until the end of experiment.
@@
Figures 15 and 15 display the maximum flow completion time (FCT) for various message sizes using STrack and RoCEv2 in 4:1 and 8:1 over-subscription networks, respectively. STrack consistently outperforms RoCEv2 by up to $3$X across all message sizes. STrack’s adaptive packet spraying mechanism effectively balances the load across inter-switch links, fully utilizing all available path bandwidth. Additionally, its congestion control reduces the window size by factors of $4$ or $8$ to accommodate the 4:1 or 8:1 over-subscription. In contrast, RoCEv2’s single path strategy suffers from heavy hash collisions, leading to network congestion, a reduction in sending rate by DCQCN, and under utilization of the network. As congestion control takes the main role to drive the performance in over-subscription scenarios, we observe that adaptive packet spraying and oblivious packet spraying exhibit similar behavior due to their use of the same congestion control algorithm.
@@
Figure 20 shows the queuing delay of last-hop ToR switch for $8$-¿$1$, $32$-¿$1$, $128$-¿$1$ incast experiments with message size of $16$MB. STrack is able to stabilize at the target queuing delay across different incast degrees, which shows the robustness of the STrack’s congestion control. Note that the stabilized point of 128-¿1 incast is is slightly higher than the target queuing delay, i.e., $8\mu$s, this is because the ideal the congestion window in this case is $1.3$ packets. Due to the rounding error, the inflight packets tend to vary between 1 or 2 packets, which causes slightly elevated queueing delay.
@@ 4.3. AI/ML Workloads
The simulated topology is a full-bisection 2-tier Clos network with all links operating at the same speed as the default. We vary the link speed to demonstrate the robustness of STrack. We only demonstrate the results for the network link speed of $400$Gbps as the trend of experiment results of different link speeds are similar. We assume one NIC maps to one GPU. Each message chunk size is $128$KB to utilize the pipeline. We primarily assess a multi-job scenario involving 64 identical collectives running concurrently on a 2048-NIC cluster, with each collective group randomly distributed throughout the cluster.
@@
Figure 24 displays the maximum collective finishing time for $64$ allreduce collectives using different allreduce algorithms with STrack and RoCEv2 on a full bisectional network. STrack with adaptive packet spraying outperforms RoCEv2 with 1 QPS_PER_CONN by up to $3$x and RoCEv2 with 4 QPS_PER_CONN by up to $27.4\%$, especially with the DoubleBinaryTree algorithm. The DoubleBinaryTree algorithm exhibits a 2:1 incast property, where the limitations of DCQCN become evident. RoCEv2 with 4 QPS_PER_CONN performs better than RoCEv2 with 1 QPS_PER_CONN across various allreduce algorithms, as the increased entropy improves performance, despite higher CPU cost. Figure 28 plots the CDF of $64$ collective finishing time of DoubleBinaryTree of Figure 24 experiment. It shows that the $64$ collectives achieves similar finishing times with STrack, while there is $26.6\%$ difference between the smallest finishing time and the maximum finishing time with RoCEv2 with 4 QPS_PER_CONN, which indicates STrack not only has better tail latency of each collective, but also better fairness compares with RoCEv2. Figure 24 shows the maximum collective finishing time for $64$ AlltoAll collectives across different levels of parallelism using STrack and RoCEv2. STrack with adaptive packet spraying outperforms RoCEv2 with $1$ QPS_PER_CONN by up to $79.2\%$ and RoCEv2 with $4$ QPS_PER_CONN by up to $18.9\%$, particularly at a parallel degree of $32$. Figure 28 shows the CDF of finishing times for 64 alltoall collectives with a parallel degree of 32. This further demonstrates that STrack offers better fairness compared to RoCEv2.
@@
Figure 24 shows the maximum collective finishing time for $64$ allreduce collectives using different allreduce algorithms with STrack and RoCEv2 on a $4$:$1$ oversubscribed network. STrack with adaptive packet spraying outperforms RoCEv2 with 1 QPS_PER_CONN by up to $4.86$X and RoCEv2 with 4 QPS_PER_CONN by up to $4.13X$, especially with the DoubleBinary Tree algorithm. We observed that the collective completion time of STrack does not significantly increase in the oversubscribed network compared to that in the full bisectional network (Figure 24). This indicates that the traffic generated by the allreduce algorithm does not fully utilize the full bisectional network. However, in this scenario, the RING and HalvingDoubling configurations of RoCEv2 with 1 QPS_PER_CONN and 4 QPS_PER_CONN exhibit worse performance than that of the full bisectional bandwidth due to its poor load balancing and congestion control mechanisms. Although RoCEv2 with 4 QPS_PER_CONN performs slightly better than RoCEv2 with 1 QPS_PER_CONN across various allreduce algorithms, it does not match STrack’s load balancing efficiency, leading to the activation of DCQCN and its associated inefficiencies becoming apparent. Figure 28 shows the CDF of finishing times for $64$ allreduce collectives with DoubleBinary Tree algorithm. Comparing to RoCEv2, STrack shows much lower collective finishing time for each collective and offers better fairness.
@@ 5.1. Congestion Signals
Contemporary programmable switches offer queuing stats at their ingress points, enabling the application of cutting payload techniques to data packets (cp, ; ndp, ). Additionally, switches can duplicate header packets and send them back to the endpoints (sfc, ; bolt, ). This approach allows for the notification of network congestion status to endpoints without experiencing ongoing congestion.
@@ 5.2. Congestion Management
There have been limited prior work focused on jointly optimizing congestion control and multipath load balancing, which is the primary objective of our system. An example of such an approach is MP-RDMA (mprdma, ), which employs packet spraying during the initial window, handles fabric and incast congestion using ECN only, which requires multiple RTTs to converge upon a congestion event. In a similar vein, Scalable Reliable Datagram (SRD (srd, )) utilizes congestion control techniques similar to BBR (bbr, ) and incorporates multipath load balancing through packet spraying. However, as a proprietary scheme, it is not clear the detailed algorithms and its loss recovery scheme.
@@ 6. Conclusion
We tackle the challenge of managing collective communication for AI/ML workload without advanced switch supports, such as packet trimming or in-network-telemetry. We propose and demonstrate that STrack, a NIC-side only and hardware-offloaded reliable transport protocol, can achieve ultra-low latency, very high bandwidth utilization and well balanced network links. By taking advantage of egress-marked ECN, STrack swiftly changing path upon early congestion detection and adaptively load balancing to utilize the overall available capacity in a multipath network. Congestion window control only comes in when a majority of the paths are congested when average RTT is elevated. With a unique error recovery mechanism, STrack also ensures rapid packet recovery within a multipath context. Our evaluation results show that highly utilized and well balanced network links are the key to ideal performance for AI/ML workloads. STrack outperforms RoCEv2 up to $6$X with synthetic workloads and by $27.4$% with collective workloads, even with the optimized RoCEv2 system setup.

### v2

@@ 3.3. STrack Reliability
Upon the arrival of each packet, a receiver sends an Selective Acknowledgment (SACK) packet if one of the following conditions are met: (1) a sufficient number of bytes have been received; (2) a packet with current expected Packet Sequence Number (EPSN) is received; (3) a probing packet is received.
@@
The receiver also records the lowest PSN received, LPSN, since the last SACK. The next SACK is sent with the bitmap segment that is associated with LPSN as retransmission timeout is always triggerred from the lowest unacknowledged PSN. Hence we need to update LPSN as early as possible to prevent the sender from timing out.
@@
Loss is inferred if the out-of-order packets counter carried at SACK exceeds a threshold. The threshold is recommended to be set to the maximum of the current congestion window’s worth of packet number and min_threshold (say 5). This constant threshold prevents spurious retransmissions when the congestion window is less than min_threshold packets. The rationale for using a threshold based on the congestion window is that if a packet does not get acknowledged within a whole window’s time, it is safe to consider it as lost. This also helps us avoid a large bitmap size at receiver/sender
@@ 4.1. Experiment Setup
We compare STrack with RoCEv2 (rocev2, ), which is widely deployed transport over Ethernet in AI/ML network. We configure the constants for STrack congestion control as suggested in  Table 1 and set the maximum number of paths as $256$ for adaptive load balancing and switch ECN mark threshold $K_{min}=25\%BDP$ and $K_{max}=75\%BDP$. Switch is operated without any advanced features enabled, and drops packets if queuing exceeds $5$ BDP. RoCEv2 utilizes DCQCN (dcqcn, ) as its primary congestion control mechanism and supports single-path operation. Additionally, it necessitates Priority Flow Control (PFC) to be enabled throughout the entire network to prevent packet loss caused by buffer overflow. We implemented shared buffer architecture for PFC to best demonstrate the performance of RoCEv2. We configured the total switch buffer size to be $256$MB for a total capacity of $51.2$Tbps and scaled the buffer size accordingly for switches with different radices. For instance, the switch buffer size is set to 128MB for a 25.6Tbps switch. We set the ECN threshold to one BDP for DCQCN and configure Lossless for RoCEv2.
@@ 4.2. Synthetic Traffic
Figure 11 samples the switch queuing delay as the simulation time progresses in the $16MB$ experiment in Figure 11 for STrack and RoCEv2. There are a total of $16K$ switch queues, $8$K queues from ToR to Spine Switch and the other $8$K queues from Spine to ToR. To avoid too many logs to be processed, we only generate logs if the queuing delay is larger than $8$us (the base RTT). Hence, if we do not see any points at one time moment in Figure 11, it means that across the $16$K queues, there is not any queue whose queue depth exceeds the BDP at that time point. Figure 11 shows that, initially before $370$us, STrack’s adaptive packet spraying experiences mild congestion due to hash collisions, but it adapts to skip the bad paths, and finds the good paths. Hence after $370$us, all the flows settle down to the good paths, and thus, the switch queuing delays do not exceed $8$us after $370$us. In contrast, RoCEv2 continues experiencing hash collisions, which leads to the queuing delay up to $250$us until the end of experiment.
@@
Figures 15 and 15 display the maximum flow completion time (FCT) for various message sizes using STrack and RoCEv2 in 4:1 and 8:1 over-subscription networks, respectively. STrack consistently outperforms RoCEv2 by up to $3$X across all message sizes. STrack’s adaptive packet spraying mechanism effectively balances the load across inter-switch links, fully utilizing all available path bandwidth. Additionally, its congestion control reduces the window size by factors of $4$ or $8$ to accommodate the 4:1 or 8:1 over-subscription. In contrast, RoCEv2’s single path strategy suffers from heavy hash collisions, leading to network congestion, a reduction in sending rate by DCQCN, and under utilization of the network. As congestion control plays the main role in driving the performance in over-subscription scenarios, we observe that adaptive packet spraying and oblivious packet spraying exhibit similar behavior due to their use of the same congestion control algorithm.
@@
Figure 20 shows the queuing delay of last-hop ToR switch for $8$-¿$1$, $32$-¿$1$, $128$-¿$1$ incast experiments with message size of $16$MB. STrack is able to stabilize at the target queuing delay across different incast degrees, which shows the robustness of the STrack’s congestion control. Note that the stabilized point of 128-¿1 incast is is slightly higher than the target queuing delay, i.e., $8\mu$s, this is because the ideal congestion window in this case is $1.3$ packets. Due to rounding error, the inflight packets tend to vary between 1 or 2 packets, which causes slightly elevated queueing delay.
@@ 4.3. AI/ML Workloads
The simulated topology is a full-bisection 2-tier Clos network with all links operating at the same speed as the default. We vary the link speed to demonstrate the robustness of STrack. We only show the results for the network link speed of $400$Gbps as the experiment results of different link speeds are similar. We assume one NIC maps to one GPU. Each message chunk size is $128$KB to utilize the pipeline. We primarily assess a multi-job scenario involving 64 identical collectives running concurrently on a 2048-NIC cluster, with each collective group randomly distributed throughout the cluster.
@@
Figure 24 displays the maximum collective finishing time for $64$ allreduce collectives using different allreduce algorithms with STrack and RoCEv2 on a full bisectional network. STrack with adaptive packet spraying outperforms RoCEv2 with 1 QPS_PER_CONN by up to $3$x and RoCEv2 with 4 QPS_PER_CONN by up to $27.4\%$, especially with the DoubleBinaryTree algorithm. The DoubleBinaryTree algorithm exhibits a 2:1 incast property, where the limitations of DCQCN become evident. RoCEv2 with 4 QPS_PER_CONN performs better than RoCEv2 with 1 QPS_PER_CONN across various allreduce algorithms, as the increased entropy improves performance, albeit with higher CPU cost. Figure 28 plots the CDF of $64$ collective finishing time of DoubleBinaryTree of Figure 24 experiment. It shows that the $64$ collectives achieves similar finishing times with STrack, while there is $26.6\%$ difference between the smallest finishing time and the maximum finishing time with RoCEv2 with 4 QPS_PER_CONN, which indicates STrack not only has better tail latency of each collective, but also better fairness compares with RoCEv2. Figure 24 shows the maximum collective finishing time for $64$ AlltoAll collectives across different levels of parallelism using STrack and RoCEv2. STrack with adaptive packet spraying outperforms RoCEv2 with $1$ QPS_PER_CONN by up to $79.2\%$ and RoCEv2 with $4$ QPS_PER_CONN by up to $18.9\%$, particularly at a parallel degree of $32$. Figure 28 shows the CDF of finishing times for 64 alltoall collectives with a parallel degree of 32. This further demonstrates that STrack offers better fairness compared to RoCEv2.
@@
Figure 24 shows the maximum collective finishing time for $64$ allreduce collectives using different allreduce algorithms with STrack and RoCEv2 on a $4$:$1$ oversubscribed network. STrack with adaptive packet spraying outperforms RoCEv2 with 1 QPS_PER_CONN by up to $4.86$X and RoCEv2 with 4 QPS_PER_CONN by up to $4.13X$, especially with the DoubleBinary Tree algorithm. We observed that the collective completion time of STrack does not significantly increase in the oversubscribed network compared to that in the full bisectional network (Figure 24). This indicates that the traffic generated by the allreduce algorithm does not fully utilize the full bisectional network. However, in this scenario, the RING and HalvingDoubling configurations of RoCEv2 with 1 QPS_PER_CONN and 4 QPS_PER_CONN exhibit worse performance than that of the full bisectional bandwidth due to its poor load balancing and congestion control mechanisms. Although RoCEv2 with 4 QPS_PER_CONN performs slightly better than RoCEv2 with 1 QPS_PER_CONN across various allreduce algorithms, it does not match STrack’s load balancing efficiency, triggering DCQCN to cut rate that leads to poor performance. Figure 28 shows the CDF of finishing times for $64$ allreduce collectives with DoubleBinary Tree algorithm. Comparing to RoCEv2, STrack shows much lower collective finishing time for each collective and offers better fairness.
@@ 5.1. Congestion Signals
Contemporary programmable switches offer queuing stats at their ingress points, enabling the application of cutting payload techniques to data packets (cp, ; ndp, ). Additionally, switches can duplicate header packets and send them back to the endpoints (sfc, ; bolt, ). This approach requires the notification of network congestion status to endpoints even without experiencing ongoing congestion.
@@ 5.2. Congestion Management
There have been limited prior work focused on jointly optimizing congestion control and multipath load balancing, which is the primary objective of our system. One example of such an approach is MP-RDMA (mprdma, ), which employs packet spraying during the initial window, and handles fabric and incast congestion using ECN only. As a result, the algorithm converges slowly as it requires multiple RTTs to reach stability upon a congestion event. In a similar vein, Scalable Reliable Datagram (SRD (srd, )) utilizes congestion control techniques similar to BBR (bbr, ) and incorporates multipath load balancing through packet spraying. However, as a proprietary scheme, it is not clear the detailed algorithms and its loss recovery scheme.
@@ 6. Conclusion
We tackle the challenge of managing collective communication for AI/ML workload without advanced switch supports, such as packet trimming or in-network-telemetry. We propose and demonstrate that STrack, a NIC-side only and hardware-offloaded reliable transport protocol, can achieve ultra-low latency, very high bandwidth utilization and well balanced networks. By taking advantage of egress-marked ECN, STrack swiftly changing path upon early congestion detection and adaptively load balancing to utilize the overall available capacity in a multipath network. Congestion window control only comes in when a majority of the paths are congested when average RTT is elevated. With a unique error recovery mechanism, STrack also ensures rapid packet recovery within a multipath context. Our evaluation results show that highly utilized and well balanced network links are the key to ideal performance for AI/ML workloads. STrack outperforms RoCEv2 up to $6$X with synthetic workloads and by $27.4$% with collective workloads, even with the optimized RoCEv2 system setup.